# A game-theoretic probability approach to loopholes in CHSH experiments

### v2

@@ A Proofs of Theorems presented in Section 4
$\leq\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}|\omega^{1}\omega^{3}|(1\pm\omega^{2}\omega^{4})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ +\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}|\omega^{1}\omega^{4}|(1\pm\omega^{2}\omega^{3})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}(1\pm\omega^{2}\omega^{4})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ +\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}(1\pm\omega^{2}\omega^{3})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{2},\omega^{4})\in\Omega^{2}}(1\pm\omega^{2}\omega^{4})\cdot\frac{S_{n}(\omega^{2},\omega^{4})}{n}+\sum_{(\omega^{2},\omega^{3})\in\Omega^{2}}(1\pm\omega^{2}\omega^{3})\cdot\frac{S_{n}(\omega^{2},\omega^{3})}{n}+o(1)$
$=\sum_{(\omega^{2},\omega^{4})\in\Omega^{2}}(1\pm\omega^{2}\omega^{4})\cdot\frac{S_{n}(\omega^{2},\omega^{4};2,4)}{T_{n}(2,4)}+\sum_{(\omega^{2},\omega^{3})\in\Omega^{2}}(1\pm\omega^{2}\omega^{3})\cdot\frac{S_{n}(\omega^{2},\omega^{3};2,3)}{T_{n}(2,3)}+o(1)$

### v3

@@ A Proofs of Theorems presented in Section 4
$\leq\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}|\omega^{1}\omega^{3}|(1\pm\omega^{2}\omega^{4})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ +\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}|\omega^{1}\omega^{4}|(1\pm\omega^{2}\omega^{3})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}(1\pm\omega^{2}\omega^{4})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ +\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}(1\pm\omega^{2}\omega^{3})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{2},\omega^{4})\in\Omega^{2}}(1\pm\omega^{2}\omega^{4})\cdot\frac{\widetilde{V_{n}}(\omega^{2},\omega^{4})}{n}+\sum_{(\omega^{2},\omega^{3})\in\Omega^{2}}(1\pm\omega^{2}\omega^{3})\cdot\frac{\widetilde{V_{n}}(\omega^{2},\omega^{3})}{n}+o(1)$
$=\sum_{(\omega^{2},\omega^{4})\in\Omega^{2}}(1\pm\omega^{2}\omega^{4})\cdot\frac{\widetilde{U_{n}}(\omega^{2},\omega^{4};2,4)}{W_{n}(2,4)}+\sum_{(\omega^{2},\omega^{3})\in\Omega^{2}}(1\pm\omega^{2}\omega^{3})\cdot\frac{\widetilde{U_{n}}(\omega^{2},\omega^{3};2,3)}{W_{n}(2,3)}+o(1)$